OpenStudy (anonymous):

Analyze the graph of f(x)=sec 2x Is there any continuity? symmetry? asymptotes? note end behavior!

OpenStudy (anonymous):

\[ \sec(2x)=\frac{1}{\cos(2x)} \]Can't have any roots, looks like it's undefined when \(2x=\pi/2 + n\pi\)